OpenStudy (anonymous):

factor 16x^2+49. check your work

OpenStudy (ranga):

a^2 + b^2 = (a + bi)(a - bi) where i is square root of -1.

OpenStudy (anonymous):

do I use that for x^2 + 64 too?

OpenStudy (ranga):

Not sure what you are referring to ^^^. 16x^2 + 49 = (4x)^2 + (7)^2 = (4x + 7i)(4x - 7i)

OpenStudy (ranga):

Oh, yes. The same formula for a different problem.

OpenStudy (anonymous):

okay, thank you! I know what you're talking about :)

OpenStudy (ranga):

You are welcome.